# A robust mixed-effects quantile regression model using generalized Laplace mixtures to handle outliers and skewness

### v1

@@ 5.4 Quantile fits
6   Simulation study
6.1 Model performance
@@
Random effects $\bm{b}_{i}=\left(b_{i1},b_{i2}\right)^{\prime}$ were drawn from a bivariate normal distribution with zero mean and covariance matrix:
@@
The model performance results are presented in Web LABEL:tab:CGAL_PERFORM of the Supplementary Material. Coverage probabilities for the model parameters range from acceptable (meeting nominal targets) to conservative, with none being overly lenient. However, higher contamination levels lead to increased bias in $\beta_{2}$, and the variance components also exhibit substantial bias, consistent with the findings by Burger et al. [23] for the covariance matrix under the MGH-$t$ prior distribution.
6.2 Robustness under data contamination
@@ 7   Discussion
An important practical consideration is that the cGAL model, while more robust to outliers than its AL and GAL counterparts, is computationally more demanding. Unlike AL and GAL, which have convenient mixture representations facilitating sampling, the cGAL distribution requires a custom likelihood formulation (via the “ones trick" in JAGS), substantially increasing runtime. This trade-off may be critical in large-scale studies or real-time clinical settings where computational efficiency is paramount. Nonetheless, the computational overhead may be justified when data contain influential outliers, given the cGAL’s superior robustness and improved predictive performance in our results.
@@
Although we treated observed viral load data in a straightforward manner, future research may account for left-censored observations below the limit of quantification (LLOQ), which is common in HIV studies. In principle, the NLME model could handle this by treating sub-threshold observations as censored. However, implementing censoring under the GAL or cGAL distributions is nontrivial due to the lack of closed-form CDFs. A specialized sampling scheme or additional numerical integration might be required to accommodate censoring within the robust quantile framework. Similarly, incorporating measurement error in the CD4 covariate could refine the modeling of its effect on long-term viral decay [24]. Presently, we assume that CD4 was measured without error, but realistic clinical data often include substantial variability. Approaches such as Bayesian measurement-error models or joint modeling of CD4 and viral load could better characterize how immune status covaries with viral trajectory.
@@ Appendix B Kullback-Leibler divergence
To evaluate the influence of individual data points on the model, we use LOO to measure model adequacy following the approach detailed by Wang and Luo [18]. Specifically, we compute the K-L divergence, quantifying the difference between model fits when a particular observation is included versus excluded.
Let $\bm{\eta}$ denote the collection of parameters. For the AL model, $\bm{\eta}=\left(\bm{\beta}^{\prime},\sigma\right)^{\prime}$ includes the fixed effects vector $\bm{\beta}$ and scale parameter $\sigma$. For the GAL model, $\bm{\eta}=\left(\bm{\beta}^{\prime},\sigma,\gamma\right)^{\prime}$, where $\gamma$ is the skewness parameter. In the cGAL model, $\bm{\eta}=\left(\bm{\beta}^{\prime},\sigma,\gamma,\alpha\right)^{\prime}$, where $\alpha$ is the contamination parameter. Let $\bm{\Theta}$ represent the complete set of model parameters, including both fixed and random effects.
@@
This K-L divergence is calculated for each observation $y_{ij}$ to evaluate its potential influence under regression model $R$, identifying observations that may substantially impact the parameter estimates. Following Tomazella et al. [19], we flag observation $y_{ij}$ as influential (“outlier") if it meets the following criterion:
@@ Acknowledgements
WEB TABLE 1: Posterior summaries for each quantile level under the AL, GAL, and cGAL models. Columns report each parameter’s posterior estimate (Est.) and 95% HPD interval. Results for higher quantiles continue on subsequent pages.

### v2

@@ Introduction
Sensitivity analyses
To confirm that our conclusions are not driven by the baseline priors, we re-estimated the cGAL model with a markedly flatter specification,
$\sigma\sim t_{+}\left(0,100,1\right),\qquad B\sim\mathrm{Beta}\left(\tfrac{1}{2},\tfrac{1}{2}\right),\qquad\alpha\sim\mathrm{Beta}\left(\tfrac{1}{2},\tfrac{1}{2}\right).$ (28)
The half-Cauchy, denoted by $t_{+}\left(0,100,1\right)$, contributes virtually no information about the scale $\sigma$. Jeffreys’ $\mathrm{Beta}\!\left(\tfrac{1}{2},\tfrac{1}{2}\right)$ places extra mass near the endpoints of the admissible intervals for $B$ (and, by transformation, $\gamma$) and $\alpha$ (Jeffreys, 1998), letting the data rather than the prior determine where those parameters lie.
We refer to the sensitivity-prior variant as $\text{cGAL}_{\text{sens}}$. Web LABEL:tab:POST_SUMMARY shows that parameter estimates and 95% HPD intervals under these flatter priors are almost unchanged from the baseline fit, confirming that our substantive conclusions are robust to prior choice.
We re-estimated the model with $\tau_{0}\in\left\{6.\dot{6}\dot{7},20,40\right\}$. Web LABEL:tab:POST_SUMMARY_CONTAM shows that all fixed-effect parameter estimates and their 95% HPD intervals are very similar to the baseline fit ($\tau_{0}=10$), confirming that our results are not sensitive to the particular choice of $\tau_{0}$.
Simulation study
Model performance
@@
Random effects $\bm{b}_{i}=\left(b_{i1},b_{i2}\right)^{\top}$ were drawn from a bivariate normal distribution with zero mean and covariance matrix:
@@
The model performance results are presented in Web LABEL:tab:CGAL_PERFORM of the Supplementary Material. Coverage probabilities for the model parameters range from acceptable (meeting nominal targets) to conservative, with none being overly lenient. However, higher contamination levels lead to increased bias in $\beta_{2}$, and the variance components also exhibit substantial bias, consistent with the findings by Burger et al. (2020) for the covariance matrix under the MGH-$t$ prior distribution.
Robustness under data contamination
@@ Discussion
An important practical consideration is that the cGAL model, while more robust to outliers than its AL and GAL counterparts, is computationally more demanding. Unlike AL and GAL, which have convenient mixture representations facilitating sampling, the cGAL distribution requires a custom likelihood formulation (via the “ones trick” in JAGS), substantially increasing runtime. This trade-off may be critical in large-scale studies or real-time clinical settings where computational efficiency is paramount. Nonetheless, the computational overhead may be justified when data contains influential outliers, given the cGAL’s superior robustness and improved predictive performance in our results.
@@
Although we treated observed viral load data in a straightforward manner, future research may account for left-censored observations below the limit of quantification (LLOQ), which is common in HIV studies. In principle, the NLME model could handle this by treating sub-threshold observations as censored. However, implementing censoring under the GAL or cGAL distributions is nontrivial due to the lack of closed-form CDFs. A specialized sampling scheme or additional numerical integration might be required to accommodate censoring within the robust quantile framework. Similarly, incorporating measurement error in the CD4 covariate could refine the modeling of its effect on long-term viral decay (Wu, 2002). Presently, we assume that CD4 was measured without error, but realistic clinical data often include substantial variability. Approaches such as Bayesian measurement-error models or joint modeling of CD4 and viral load could better characterize how immune status covaries with viral trajectory.
@@ Appendix B Kullback-Leibler divergence
To evaluate the influence of individual data points on the model, we use LOO to measure model adequacy following the approach detailed by Wang and Luo (2016). Specifically, we compute the K-L divergence, quantifying the difference between model fits when a particular observation is included versus excluded.
Let $\bm{\eta}$ denote the collection of parameters. For the AL model, $\bm{\eta}=\left(\bm{\beta}^{\top},\sigma\right)^{\top}$ includes the fixed effects vector $\bm{\beta}$ and scale parameter $\sigma$. For the GAL model, $\bm{\eta}=\left(\bm{\beta}^{\top},\sigma,\gamma\right)^{\top}$, where $\gamma$ is the skewness parameter. In the cGAL model, $\bm{\eta}=\left(\bm{\beta}^{\top},\sigma,\gamma,\alpha\right)^{\top}$, where $\alpha$ is the contamination parameter. Let $\bm{\Theta}$ represent the complete set of model parameters, including both fixed and random effects.
@@
This K-L divergence is calculated for each observation $y_{ij}$ to evaluate its potential influence under regression model $R$, identifying observations that may substantially impact the parameter estimates. Following Tomazella et al. (2021), we flag observation $y_{ij}$ as influential (“outlier”) if it meets the following criterion:
@@ Acknowledgements
[…]